QUESTION 2 Administer 125 mL in 45 minutes using a microdrip set. Caleulate the gtt/min to the nearest whole number. O A 166 gttlmin O B 166.7 gttmin O C 167gtt/min O D 168gttmin QUESTION 3 The order is to infuse 1300 mL in 8 hours using a 20 gtt/mL set. Calculate the gttmin to the nearest whole number O A. 54 gtt/min O B. 162.5 gt/min ° C. 54.1 gtt/min Infuse 50 mL in 15 minutes using a 10 gt/mlL set. Calculate the gtt/min to the nearest whole number O A 33.3 gt/min O B. 3.3 gtt/min QUESTION 4 O D 3.3 mL/min
Answer #1

2. Flow rate = volume/time x drop factor

Microdrip set have a drop factor of 60gtt/ml

Flow rate = 125/45 x 60 = 166.7gtt/min

The nearest whole number is 167gtt/min

So the answer is (C)

3. Flow rate = volume(ml)/ time(min) x drop factor

Flow rate =( 1300/ 8x60) x 20 = 54.1gtt/min

So the nearest whole number is 54gtt/min

So the answer is (A)

4. Flow rate = 50/15 x 10 = 33.3gtt/min

The nearest whole number is 33gtt/min

So the answer is (c)